OpenStudy (maheshmeghwal9):

If x=p+2, find p from the following equation.

OpenStudy (maheshmeghwal9):

\[p-(x-3/3)=2/3(4-x)\]

OpenStudy (anonymous):

Multiply by 3 on both LHS and RHS. Then substitute p = x-2 and find x.

OpenStudy (lgbasallote):

since x = p + 2 substute p+2 to every x p - (p+2 - 3/3) = 2/3(4 - p - 2)

OpenStudy (anonymous):

After you find x, substitute your x in p = x-2 to find p

OpenStudy (maheshmeghwal9):

ok!